89988989:

Maria's family drove 140 mi to her grandparents' house and averaged 56 mi/h​ on the way there. On the return trip, they averaged 50 mi/h. What was the total time Maria's family spent driving to and from her grandparents' house?

Shadow:

\[(140 mi \times \frac{ h }{ 56mi }) + (140mi \times \frac{ h }{ 50mi }) = totalhours\]

Shadow:

140 miles is the distance. Multiply that by the velocity (basically speed) and you get the time spent traveling. You just set up the fractions so that miles is cancelling out and hours remains.

Shadow:

Basically you should be adding 140/56 with 140/50.